OpenStudy (anonymous):

what is the value of w in the equation 0.04w+0.6=2.4

OpenStudy (anonymous):

Alright, so first you have to leave w by itself on one side. So what you do is move 0.6 to the other side. 0.04w=2.4-0.6=1.8 Then divide 0.04 to the other side w=1.8/0.04 You can solve from there.